OpenStudy (anonymous):

(5.4x10 to the neg 1)(7x10 to the 0 power)

OpenStudy (lalaly):

\[(5.4 \times 10^{-1})(7 \times 10^0)\]\[=\frac{5.4}{10} \times 7\]\[=3.78\]

OpenStudy (anonymous):

(5.4*10^-1)*(7*10^0) =(5.4/10)*(7*1) =0.54*7 =3.78